OpenStudy (anonymous):

Given QRS equals to TUV, QS = 3v+2 and TV = 7v – 6, find the length of QS and TV.

OpenStudy (anonymous):

what type of question is this????

OpenStudy (anonymous):

a Geometry question

OpenStudy (anonymous):

What's QRS? A triangle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In that case then 3v+2 = 7v-6?

OpenStudy (anonymous):

okay but then after that i get 2+4-6 and i add 6 but what i am confused about is would my answers be 8 or 4v ???

OpenStudy (anonymous):

3v+2 = 7v-6 so 4v = 8 and v= 2 QS = 3v+2 = 8 and TV = 7v-6 = 8 as well

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

ur welcome